OpenStudy (anonymous):

a company has 350 workers. the president of the company wants to know what percent increase in employmentwould be necessary for the number of workers to be greater than 375. help me please

OpenStudy (triciaal):

number of workers plus increase now number of workers more than 375 increase percentage let's use x 350 + x%of 350 > 375 first, what % will give 375 x/100 (350) =25 means x = 7.142857 as a factor increases the product increases 7%of 350 =25 350 +7%(350) = 375 > 7% of 350 means > 25 350 + (> 7% of 350 ) = 350 + (> 25) = >375 assuming we are only using whole number for the percentage increase then it would be 8% minimum increase to have >375 workers